# Masked Language Models are Good Heterogeneous Graph Generalizers

### v1

@@ Abstract
Heterogeneous graph neural networks (HGNNs) excel at capturing structural and semantic information in heterogeneous graphs (HGs), while struggling to generalize across domains and tasks. Recently, some researchers have turned to integrating HGNNs with large language models (LLMs) for more generalizable heterogeneous graph learning. However, these approaches typically extract structural information via HGNNs as HG tokens, and disparities in embedding spaces between HGNNs and LLMs have been shown to bias the LLM’s comprehension of HGs. Moreover, as these HG tokens are often derived from node-level tasks, the model’s ability to generalize across tasks remains limited. To this end, we propose a simple yet effective Masked Language Modeling-based method, called MLM4HG. MLM4HG introduces metapath-based textual sequences instead of HG tokens to extract structural and semantic information inherent in HGs, and designs customized textual templates to unify different graph tasks into a coherent cloze-style <mask> token prediction paradigm. Specifically, MLM4HG first converts HGs from various domains to texts based on metapaths, and subsequently combines them with the unified task texts to form a HG-based corpus. Moreover, the corpus is fed into a pretrained LM for fine-tuning with a constrained target vocabulary, enabling the fine-tuned LM to generalize to unseen target HGs. Extensive cross-domain and multi-task experiments on four real-world datasets demonstrate the superior generalization performance of MLM4HG over state-of-the-art methods in both few-shot and zero-shot scenarios. Our code is available at https://github.com/BUPT-GAMMA/MLM4HG.
@@ 1 Introduction
To validate our model’s generalization ability, we conduct extensive experiments on four real-world datasets from distinct domains and graph tasks in both few-shot and zero-shot scenarios. Our experiments follow the leave-one-out strategy Ding et al. (2023), where the model is fine-tuned on source HGs across different graph tasks and then adapted to an unseen target HG. The results demonstrate that MLM4HG outperforms the best-performing baselines by an average of 18.08% in Micro-F1 and 7.10% in Macro-F1 in few-shot settings, as well as 15.22% in Micro-F1 and 8.04% in Macro-F1 in zero-shot settings. Our contributions are summarized as follows:
@@ 2.1 Heterogeneous Graph Representation Learning
Heterogeneous graph neural networks (HGNNs) Wang et al. (2022); Yang et al. (2021) and self-supervised learning (SSL) Yang et al. (2022) are effective methods for heterogeneous graph representation learning. Existing HGNNs can be broadly categorized into relation-based methods Hu et al. (2020b); Lu et al. (2019); Schlichtkrull et al. (2018), which capture heterogeneous structural information by treating different edge types as distinct relations, and metapath-based methods Dong et al. (2017); Fu et al. (2020); Wang et al. (2019b), which consider the semantic information inherent in HGs. For example, RHINE Lu et al. (2019) distinguish the various relations into Affiliation Relations (ARs) with one-centered by-another structures and Interaction Relations (IRs) with peer-to-peer structures. HAN Wang et al. (2019b) adopts hierarchical attention to aggregate information at both the node and semantic levels. SSL methods include contrastive methods and generative methods. Contrastive methods Jiang et al. (2021); Wang et al. (2021) aim to maximize mutual information between different views. For example, HeCo Wang et al. (2021) employs network schema and meta-path views to assist the model in learning high-level node embeddings. Generative methods Hu et al. (2020a); Tian et al. (2023), on the other hand, leverage graph reconstruction and property prediction to extract general structural and attribute semantics. However, these methods focus on a single HG domain or a specific task, lacking the capacity to generalize across diverse HGs and tasks.
@@ 4.2 Masked Language Modeling for Generalizable HG Learning
Why Masked LMs Instead of LLMs? MLM4HG reformulates various tasks into a unified cloze-style <mask> token prediction paradigm, which essentially corresponds to classification. Prior work Zhang et al. (2024) shows that, compared to causal language modeling-based LLMs, masked LMs are better suited for such tasks both empirically and theoretically. Consequently, masked LMs align more naturally with the cloze-style formulation. Empirically, they achieve superior performance (Sections 5.2 and 5.3) while incurring significantly lower computational overhead (Appendix E).
@@ 5 Experiments and Analysis
RQ4: Do metapath-based sequences extract structural information for downstream tasks?
RQ5: How effective are metapath-based sequences compared to other graph-to-text methods?
Additionally, we further investigate the impact of using different LMs or LLMs as backbones in Appendix E. We also analyze the influence of various hyperparameter settings on the experimental results, as presented in Appendix F. To provide an intuitive understanding of the model’s behavior, we conduct a visual analysis of the attention scores from the last layer of MLM4HG across different graph tasks in Appendix H.
@@ 5.1 Experimental Settings
Datasets. We conduct extensive experiments on four real-world benchmark datasets from different domains: IMDB, DBLP Wang et al. (2019b), YELP, and PubMed Yang et al. (2020). More details can be found in Appendix A.
Baselines. To comprehensively evaluate the performance of our MLM4HG, we compare it with the following five representative and state-of-the-art methods from three different categories. The first category is represented by the HGNN method HAN Wang et al. (2019b). The second category is the heterogeneous graph self-supervised learning method HeCo Wang et al. (2021). The third category includes LM and LLM-enhanced graph methods, including WalkLM Tan et al. (2024), GraphGPT Tang et al. (2024a), and HiGPT Tang et al. (2024b).
@@
Implementation Details. The LM used in our experiments is DistilRoBERTa Sanh (2019). To automatically generate diverse and meaningful metapaths while reducing reliance on manual design, we comprehensively traverse all edge types to construct candidate metapaths, as shown in Appendix A. For each node in the HG, we randomly sample three instances per metapath. This design choice is motivated by our observation (Appendix F) that longer metapath-based sequences enable the LM to better capture intricate structural and semantic information. Notably, sampling three instances yields sequence lengths that closely align with the model’s maximum input capacity of 512 tokens, thereby fully utilizing its contextual modeling capabilities. The LM is fine-tuned for three epochs, and the other hyperparameters of MLM4HG are optimized through grid-searching for best performance. The baseline parameters are initially set according to the values reported in the original papers and then optimized through grid-searching to achieve optimal performance. All hyperparameter settings are provided in Appendix C, and a detailed hyperparameter study is available in Appendix F.
@@ 5.2 Cross-domain Generalization Analysis (RQ1)
The overall experimental results are presented in Table 1. (1) MLM4HG outperforms various state-of-the-art baselines across 16 groups in four datasets, with an average improvement of 18.08% in Micro-F1 and 7.10% in Macro-F1. (2) MLM4HG with 1-shot outperforms other methods with 40-shot in IMDB and PubMed datasets, demonstrating that MLM4HG can capture general knowledge from source HGs and effectively generalize to the target HG. (3) HiGPT surpasses GraphGPT and demonstrates the second-best performance in most datasets, as it adopts heterogeneous graph instruction tuning to achieve heterogeneous relation awareness. However, the HGNN-generated HG tokens introduce bias into the LLM’s understanding, and the ignorance of metapaths further leading to suboptimal results. (4) WalkLM outperforms HiGPT on the PubMed dataset, likely because the random walk-based textual sequences in WalkLM allow the LM to directly comprehend the HG. In contrast, HGNN-generated HG tokens in HiGPT introduce additional complexity for the LLM when handling PubMed’s specialized medical content. However, WalkLM fails to account for generalization and omits metapaths, which ultimately leads to suboptimal performance.
5.3 Multi-task Generalization Analysis (RQ2)
@@
5.4 Ablation Study (RQ3)
In this subsection, we introduce four variants to examine the influence of different learning scenarios and model components on the experimental results. To capture more comprehensive and general graph knowledge from source HGs, MLM4HG adopts cross-domain and multi-task fine-tuning by default. To evaluate whether the model benefits from this learning scenarios, we define two variants: w/o Fine-tuning and w/o Multi-task. w/o Fine-tuning indicates that the model is directly applied to the target HG without cross-domain and multi-task fine-tuning on source HGs. w/o Multi-task refers to fine-tuning the pretrained LM on source HGs using the single same task as the target HG. Additionally, to assess the impact of different components on the experimental outcomes, we design two variants: w/o Metapath and w/o Constraint. In w/o Metapath, the metapath-based sequences is replaced with node attributes. In w/o Constraint, the predicted <mask> token is selected from the full vocabulary rather than the constrained target vocabulary.
We evaluated the performance of these variants in node classification and link prediction tasks in both few-shot and zero-shot scenarios. The experimental results are shown in Table 3. (1) It is evident that these variants perform suboptimally compared to the original MLM4HG across different graph tasks and experimental settings. (2) w/o Multi-task outperforms w/o Fine-tuning as even single-task fine-tuning on source HGs enables the LM to extract transferable information, significantly improving performance on the target dataset. Nonetheless, the incorporation of multi-task fine-tuning in MLM4HG enables the LM to extract richer information from multiple perspectives, further enhancing downstream inference. (3) When the vocabulary size is unconstrained in w/o Constraint, the results experience a significant decline compared to MLM4HG. This clearly indicates that the model’s focus should be directed toward the target categories, especially for our simpler graph tasks. (4) When node attributes replace metapath-based sequences in w/o Metapath, the performance of both node classification and link prediction tasks also decreases to varying degrees. This confirms that metapath-based sequences contains rich structural and semantic information, which helps the LM extract general heterogeneous graph knowledge and generalizing across different HGs.
5.5 Evaluating Structural Information in Metapath-based Sequences (RQ4)
To verify whether the LM can benefit from both structural and semantic information extracted from the metapath-based sequences, we design and implement an exploratory experiment focused on the supervised node classification task using the IMDB dataset. As shown in Figure 2, we process the metapath-based sequences into three types: Attribute Only, Structure Only, and Both Attribute and Structure. In Attribute Only, only attribute information is included. In Structure Only, all nodes are anonymized with no attribute information, while in Both Attribute and Structure, both structure and attribute information are provided. The results are shown in Figure 2. In the 20% supervised learning scenarios, all three input types outperform the metapath-based HGNN baseline, with Both Attribute and Structure achieving the best results. This highlights the effectiveness of metapath-based sequences in allowing the LM to integrate both structural and semantic information from various source heterogeneous graphs. While the LM’s prior knowledge contributes to the strong performance of Attribute Only, the competitive results of Structure Only indicate that our approach does not merely rely on pretrained knowledge. Instead, the LM is able to extract and leverage structural signals from metapath-based sequences for downstream tasks.
5.6 Graph-to-Text Method Comparison (RQ5)
We compare our metapath-based sequences with other commonly used graph-to-text methods. Specifically, we compare the GML Himsolt (1997) and GraphML Brandes et al. (2013) methods with our metapath-based sequences. Networkx provides a unified parser for these two formats, thus we use the parsed node and edge information as a replacement for the metapath-based sequences. Examples and parsing results of these two formats are provided in Appendix I.
As shown in Table 4, the metapath-based sequences demonstrate superior performance in both node classification and link prediction tasks on the IMDB and YELP datasets. Compared to our metapath-based sequences, GML and GraphML merely offer a straightforward description of nodes and edges, requiring the LM to further extract the rich and implicit semantic information, which increases the task complexity. This confirms that our metapath-based sequences effectively extracts both structural and semantic information from HGs, benefiting downstream tasks.
@@ 6 Conclusion
In this paper, we propose a simple yet effective Masked Language Modeling-based approach for generalizable Heterogeneous Graph learning, called MLM4HG. MLM4HG introduces metapath-based sequences to unify the representation of HGs, enabling cross-domain generalization. Then it reformulates different tasks into a cloze-style <mask> token prediction paradigm using customized templates, thereby supporting multi-task generalization. By constructing an HG-based corpus and using it to fine-tune a pretrained LM, we enable the LM to generalize across unseen target graphs and tasks. To the best of our knowledge, we are the first to apply masked language models for generalizable heterogeneous graph learning, establishing a new paradigm and paving a new path for future research. Extensive experiments in four real-world datasets from distinct domains demonstrate MLM4HG’s superior performance across different graph tasks in both few-shot and zero-shot scenarios.
Limitations. While MLM4HG effectively captures structural signals from metapath-based sequences (Section 5.5), the semantic information from textual content remains crucial for LMs. Thus, applying LMs to non-textual graph data still requires further exploration. Moreover, additional investigation is needed to assess the model’s generalizability across a wider range of tasks, such as regression tasks.
Broader Impacts. MLM4HG holds promise for robust AI development in real-world applications involving HGs, including social networks, knowledge graphs, and recommender systems.
@@ Appendix A Dataset Statistics
IMDB Movie (M), 4,278 Actor (A), 5,257 Director (D), 2,081 acted in, 12,828 was acted by, 12,828 directed, 4,278 was directed by, 4,278 Action Comedy Drama MAM, MDM AMA, AMDMA DMD, DMAMD
DBLP Author (A), 4,057 Paper (P), 14,328 Term (T), 7,723 Conference (C), 20 write, 19,645 was written by, 19,645 publish, 85,810 was published in, 85,810 receive, 14,328 was received by, 14,328 1, 2, 3, 4 APA, APCPA, APTPA PAP, PTP, PCP TPT, TPCPT, TPAPT CPC, CPAPC, CPTPC
YELP Business (B), 7,474 Location (L), 39 Phrase (P), 74,943 Stars (S), 9 rate, 7,474 score, 7,474 indicate, 2,654,313 located in, 7,474 described with, 2,654,313 is associated with, 7,474 beauty & spas pizza, food sandwiches nightlife burgers Mexican American (new) shopping automotive Italian, bars restaurants breakfast & brunch American (traditional) event planning & services BLB, BSB, BLB LBL, LBPBL, LBSBL SBS, SBLBS, SBPBS PBP, PBLBP, PBSBP
PubMed Disease (D), 20,163 Gene (G), 13,562 Chemical (C), 26,522 Species (S), 2,863 in, 97,297 with, 8,400 contain, 88,897 causing, 25,962 is caused by, 25,962 cardiovascular disease glandular disease nervous disorder communicable disease inflammatory disease pycnosis skin disease cancer DGD, DCD, DSD GDG, GCG, GSG CDC, CGC, CSC SDS, SGS, SCS

### v2

@@ Abstract
Heterogeneous graph neural networks (HGNNs) excel at capturing structural and semantic information in heterogeneous graphs (HGs), while struggling to generalize across domains and tasks. With the rapid advancement of large language models (LLMs), a recent study explored the integration of HGNNs with LLMs for generalizable heterogeneous graph learning. However, this approach typically encodes structural information as HG tokens using HGNNs, and disparities in embedding spaces between HGNNs and LLMs have been shown to bias the LLM’s comprehension of HGs. Moreover, since these HG tokens are often derived from node-level tasks, the model’s ability to generalize across tasks remains limited. To this end, we propose a simple yet effective Masked Language Modeling-based method, called MLM4HG. MLM4HG introduces metapath-based textual sequences instead of HG tokens to extract structural and semantic information inherent in HGs, and designs customized textual templates to unify different graph tasks into a coherent cloze-style <mask> token prediction paradigm. Specifically, MLM4HG first converts HGs from various domains to texts based on metapaths, and subsequently combines them with the unified task texts to form a HG-based corpus. Moreover, the corpus is fed into a pretrained LM for fine-tuning with a constrained target vocabulary, enabling the fine-tuned LM to generalize to unseen target HGs. Extensive cross-domain and multi-task experiments on four real-world datasets demonstrate the superior generalization performance of MLM4HG over state-of-the-art methods in both few-shot and zero-shot scenarios. Our code is available at https://github.com/BUPT-GAMMA/MLM4HG.
@@ 1 Introduction
We conduct extensive cross-domain and multi-task experiments by fine-tuning the model on source HGs across different graph tasks and adapting it to an unseen target HG. The evaluation is performed on four real-world datasets from distinct domains and graph tasks in both few-shot and zero-shot scenarios. The results demonstrate that MLM4HG outperforms the best-performing baselines by an average of 18.08% in Micro-F1 and 7.10% in Macro-F1 in few-shot settings, as well as 15.22% in Micro-F1 and 8.04% in Macro-F1 in zero-shot settings. Our contributions are summarized as follows:
@@ 2.1 Heterogeneous Graph Representation Learning
Heterogeneous graph neural networks (HGNNs) Wang et al. (2022); Yang et al. (2021) and self-supervised learning (SSL) Yang et al. (2022) are effective methods for heterogeneous graph representation learning.
Existing HGNNs can be broadly categorized into two types. Relation-based methods Hu et al. (2020b); Lu et al. (2019); Schlichtkrull et al. (2018) capture heterogeneous structural information by treating different edge types as distinct relations. For example, RHINE Lu et al. (2019) distinguish the various relations into Affiliation Relations (ARs) with one-centered by-another structures and Interaction Relations (IRs) with peer-to-peer structures. Considering the semantic information inherent in HGs, lots of metapath-based methods Dong et al. (2017); Fu et al. (2020); Wang et al. (2019b) have emerged in recent years. For instance, HAN Wang et al. (2019b) adopts hierarchical attention to aggregate information at both the node and semantic levels.
SSL methods include contrastive methods and generative methods. Contrastive methods Jiang et al. (2021); Wang et al. (2021) aim to maximize mutual information between different views. For example, HeCo Wang et al. (2021) employs network schema and meta-path views to assist the model in learning high-level node embeddings. Generative methods Hu et al. (2020a); Tian et al. (2023), on the other hand, leverage graph reconstruction and property prediction to extract general structural and attribute semantics. For instance, GPT-GNN Hu et al. (2020a) utilizes self-supervised attribute generation and edge generation tasks to capture the inherent dependencies between node attributes and graph structure. However, these methods focus on a single HG domain or a specific task, lacking the capacity to generalize across diverse HGs and tasks.
@@ 4.2 Masked Language Modeling for Generalizable HG Learning
4.3 Discussion
Why Masked LMs Instead of LLMs? MLM4HG reformulates various tasks into a unified cloze-style <mask> token prediction paradigm, which essentially corresponds to classification. Prior work Zhang et al. (2024) shows that, compared to causal language modeling-based LLMs, masked LMs are better suited for such tasks both theoretically and empirically. Consequently, masked LMs align more naturally with the cloze-style formulation. Empirically, they achieve superior performance (Sections 5.2 and 5.3) while incurring significantly lower computational overhead (Appendix E).
Relying Solely on Pretrained LM’s Prior Knowledge? As discussed in this subsection, we adopt a pretrained language model with inherent prior knowledge as the backbone. To access whether the LM can effectively capture structural information embedded in metapath-based sequences, which is independent of its original knowledge, and generalize to the target dataset, we conduct an exploratory experiment on the supervised node classification task using the IMDB dataset.
As shown in Figure 2, we process the metapath-based sequences into three types: Attribute Only, Structure Only, and Both Attribute and Structure. In Attribute Only, only attribute information is included. In Structure Only, all nodes are anonymized with no attribute information, while in Both Attribute and Structure, both structure and attribute information are provided. The results are shown in Figure 2. In the 20% supervised learning scenarios, all three input types outperform the metapath-based HGNN baseline HAN Wang et al. (2019b), with Both Attribute and Structure achieving the best results. This highlights the effectiveness of metapath-based sequences in allowing the LM to integrate both structural and semantic information from various source heterogeneous graphs. While the LM’s prior knowledge contributes to the strong performance of Attribute Only, the competitive results of Structure Only indicate that our approach does not merely rely on pretrained knowledge. Instead, the LM is able to extract and leverage structural signals from metapath-based sequences for downstream tasks.
@@ 5 Experiments and Analysis
RQ4: How effective are metapath-based sequences compared to other graph-to-text methods?
Additionally, we further investigate the impact of using different LMs or LLMs as backbones in Appendix E. We also analyze the impact of different hyperparameters (Appendix F) and provide a case study (Appendix I) with a visual analysis of attention scores from the final layer of MLM4HG across various graph tasks.
@@ 5.1 Experimental Settings
Datasets. We conduct extensive experiments on four real-world heterogeneous graph benchmark datasets from different domains: IMDB, DBLP Wang et al. (2019b), YELP, and PubMed Yang et al. (2020). More details can be found in Appendix A.
Baselines. To comprehensively evaluate the performance of our MLM4HG, we compare it with the following five representative and state-of-the-art methods from three different categories. The first category is represented by the HGNN method HAN Wang et al. (2019b). The second category is the heterogeneous graph self-supervised learning method HeCo Wang et al. (2021). The third category includes LM and LLM-enhanced graph methods, including WalkLM Tan et al. (2024), GraphGPT Tang et al. (2024a), and HiGPT Tang et al. (2024b), in which HiGPT is specifically designed for heterogeneous graphs, while the remaining methods are tailored for homogeneous graphs.
@@
Implementation Details. The LM used in our experiments is pretrained DistilRoBERTa Sanh (2019). To automatically generate diverse metapaths while reducing reliance on manual design, we comprehensively traverse all edge types to construct candidate metapaths, as shown in Appendix A. For each node in HGs, we randomly sample 3 instances per metapath. This design choice is motivated by our observation (Appendix F) that longer metapath-based sequences enable the LM to better capture intricate structural and semantic information. Notably, sampling three instances yields sequence lengths that closely align with the model’s maximum input capacity of 512 tokens, thereby fully utilizing its contextual modeling capabilities. The LM is fine-tuned for three epochs, and the other hyperparameters of MLM4HG are optimized through grid-searching for best performance. The baseline parameters are initially set according to the values reported in the original papers and then optimized through grid-searching to achieve optimal performance. All hyperparameter settings are provided in Appendix C, and a detailed hyperparameter study is available in Appendix F.
@@ 5.2 Cross-domain Generalization Analysis (RQ1)
The overall experimental results are presented in Table 1. (1) MLM4HG consistently outperforms a range of state-of-the-art baselines across 16 evaluation groups on four datasets, achieving an average gain of 18.08% in Micro-F1 and 7.10% in Macro-F1. (2) Notably, MLM4HG with only 1-shot supervision surpasses other methods with 40-shot settings on the IMDB and PubMed datasets, indicating its strong ability to capture transferable knowledge from source HGs and effectively generalize to target HGs with minimal supervision. (3) HiGPT achieves the second-best performance in most cases and outperforms GraphGPT, benefitting from its heterogeneous graph instruction tuning, which helps it recognize relation types more effectively. However, the reliance on HG tokens generated by HGNNs introduces embedding space mismatch, and the lack of metapath-level semantics further hinders its performance. (4) WalkLM performs better than HiGPT on PubMed, likely because its random walk-based sequences enable the LM to more directly capture graph structures, which is especially useful for PubMed’s complex biomedical data. However, WalkLM fails to account for generalization and omits metapaths, which ultimately leads to suboptimal performance.
5.3 Multi-task Generalization Analysis (RQ2)
@@
5.4 Ablation Study (RQ3)
In this subsection, we introduce four variants to examine the influence of different learning scenarios and model components on the experimental results. To capture more comprehensive and general graph knowledge from source HGs, MLM4HG adopts cross-domain and multi-task fine-tuning by default. To evaluate whether the model benefits from this learning scenarios, we define two variants: w/o Fine-tuning and w/o Multi-task. w/o Fine-tuning indicates that the model is directly applied to the target HG without cross-domain and multi-task fine-tuning on source HGs. w/o Multi-task refers to fine-tuning the pretrained LM on source HGs using the single same task as the target HG. That is, w/o Multi-task adds multi-domain pretraining to w/o Fine-tuning, whereas the complete MLM4HG includes both multi-domain and multi-task pretraining. Additionally, to assess the impact of different components on the experimental outcomes, we design two variants: w/o Metapath and w/o Constraint. In w/o Metapath, the metapath-based sequences is replaced with node attributes. In w/o Constraint, the predicted <mask> token is selected from the full vocabulary rather than the constrained target vocabulary.
We evaluated the performance of these variants in node classification and link prediction tasks in both few-shot and zero-shot scenarios. The experimental results are shown in Table 3. (1) It is evident that these variants perform suboptimally compared to the original MLM4HG across different graph tasks and experimental settings. (2) w/o Multi-task outperforms w/o Fine-tuning and achieves competitive results overall, demonstrating that multi-domain pretraining enables the LM to extract transferable knowledge and significantly boosts performance, even when fine-tuned on a single task. Further performance gains achieved by MLM4HG highlight the benefit of multi-task fine-tuning, which facilitates the extraction of more comprehensive information from multiple perspectives. (3) When the vocabulary size is unconstrained in w/o Constraint, the results experience a significant decline compared to MLM4HG. This clearly indicates that the model’s focus should be directed toward the target categories, especially for our simpler graph tasks. (4) When node attributes replace metapath-based sequences in w/o Metapath, the performance of both node classification and link prediction tasks also decreases to varying degrees. This confirms that metapath-based sequences contains rich structural and semantic information, which helps the LM extract general heterogeneous graph knowledge and generalizing across different HGs. (5) We further observe that in the few-shot scenarios, all variants—except for w/o Constraint—perform comparably to MLM4HG. This is because the pretrained LM can quickly adapt to a small number of labeled examples, as shown in Table 1. For instance, our methods already outperform the strongest baselines under the 5-shot setting on the IMDB and YELP datasets. However, in the more challenging zero-shot scenario, each component becomes essential.
@@ 5.4 Ablation Study (RQ3)
5.5 Graph-to-Text Method Comparison (RQ4)
We compare our metapath-based sequences with commonly used graph-to-text methods, specifically GML Himsolt (1997) and GraphML Brandes et al. (2013). Networkx provides a unified parser for these two formats, thus we use the parsed node and edge information as a replacement for the metapath-based sequences. Examples and parsing results of these two formats are provided in Appendix J.
As shown in Table 4, metapath-based sequences demonstrate superior performance in both node classification and link prediction tasks on the IMDB and YELP datasets. Compared to metapath-based sequences, GML and GraphML merely offer a straightforward description of nodes and edges, requiring the LM to further extract the rich and implicit semantic information, which increases the task complexity. This confirms that our metapath-based sequences effectively extracts both structural and semantic information from HGs, benefiting downstream tasks.
@@ 6 Conclusion
In this paper, we propose a simple yet effective Masked Language Modeling-based approach for generalizable Heterogeneous Graph learning, called MLM4HG. MLM4HG introduces metapath-based sequences to unify the representation of HGs, enabling cross-domain generalization. Then it reformulates different tasks into a cloze-style <mask> token prediction paradigm using customized templates, thereby supporting multi-task generalization. By constructing an HG-based corpus and using it to fine-tune a pretrained LM, we enable the LM to generalize across unseen target graphs and tasks. To the best of our knowledge, we are the first to apply masked language models for generalizable heterogeneous graph learning, establishing a new paradigm and paving a new path for future research. Extensive experiments in four real-world datasets from distinct domains demonstrate MLM4HG’s superior performance across different graph tasks in both few-shot and zero-shot scenarios. Limitations and broader impacts are discussed in Appendix H.
@@ Appendix A Dataset Statistics
IMDB
Movie (M), 4,278
Actor (A), 5,257
Director (D), 2,081
acted in, 12,828
was acted by, 12,828
directed, 4,278
was directed by, 4,278
Action
Comedy
Drama
MAM, MDM
AMA, AMDMA
DMD, DMAMD
DBLP
Author (A), 4,057
Paper (P), 14,328
Term (T), 7,723
Conference (C), 20
write, 19,645
was written by, 19,645
publish, 85,810
was published in, 85,810
receive, 14,328
was received by, 14,328
1, 2, 3, 4
APA, APCPA, APTPA
PAP, PTP, PCP
TPT, TPCPT, TPAPT
CPC, CPAPC, CPTPC
YELP
Business (B), 7,474
Location (L), 39
Phrase (P), 74,943
Stars (S), 9
rate, 7,474
score, 7,474
indicate, 2,654,313
located in, 7,474
described with, 2,654,313
is associated with, 7,474
beauty & spas
pizza, food
sandwiches
nightlife
burgers
Mexican
American (new)
shopping
automotive
Italian, bars
restaurants
breakfast & brunch
American (traditional)
event planning & services
BLB, BSB, BLB
LBL, LBPBL, LBSBL
SBS, SBLBS, SBPBS
PBP, PBLBP, PBSBP
PubMed
Disease (D), 20,163
Gene (G), 13,562
Chemical (C), 26,522
Species (S), 2,863
in, 97,297
with, 8,400
contain, 88,897
causing, 25,962
is caused by, 25,962
cardiovascular disease
glandular disease
nervous disorder
communicable disease
inflammatory disease
pycnosis
skin disease
cancer
DGD, DCD, DSD
GDG, GCG, GSG
CDC, CGC, CSC
SDS, SGS, SCS